OpenStudy (anonymous):

Expand the following using either the Binomial Theorem or Pascal’s Triangle. You must show your work for credit. (x - 5)^5

OpenStudy (anonymous):

do you know what pascals triangle is?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I realize that I have to go to the fifth row, but don't understand what to do afterwards

OpenStudy (anonymous):

1 5 10 10 5 1

OpenStudy (anonymous):

Right! So pascals triangle essentially gives you the coefficients for an expanded perfect power polynomial So really quickly its: 1 1 2 1 1 3 3 1 1 4 6 4 1 1 5 10 10 5 1

OpenStudy (anonymous):

Okay, I get that

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

@LifeEngineer

OpenStudy (anonymous):

forgot a row btw... should be 1, 11, 121...*** so for example, a perfect cube such as (a+b)^3 could be expanded using the fourth row (the first row is for the zeroth power) of pascals triangle (a+b)^3 = 1*a^3 + 3*a^2*b + 3*a*b^2 + 1*b^3

OpenStudy (anonymous):

where pascals triangle gives you the coefficients on each term, and you increment the power on the 'a' term, and decrement the power on the 'b' term

OpenStudy (anonymous):

this makes things a good bit easier than actually expanding the whole term

OpenStudy (anonymous):

You said that the coefficients are used for Pascals triangle

OpenStudy (anonymous):

now in your case, you have (x-5)^5, so you'd need to use the 6th row of pascals triangle (1 5 10 10 5 1) to get the coefficients of each term, then go through and add your a's and b's (in this case x and -5) to the appropriate powers start with x^5*(-5)^0 which is your highest degree term. the first number in pascals triangle for the 6th row is 1 so you get 1*x^5 then the second term would be x^4 * (-5)^1, with a coefficient of 5 this gives you 5*(x^4)*(-5) = -25x^4 continuing on with the third term ( x^3*(-5)^2 ), the coefficient is 10 so your term would be 10*(x^3)*(-5)^2 and so on and so forth until all your terms are satisfied! I hope that helped, I realize my explanation wasnt the best...